OpenStudy (amtran_bus):

If xy^2 = 12 and xy = 4, what is x?

ganeshie8 (ganeshie8):

xy = 4 gives y = 4/x Plug this in the first equation

OpenStudy (amtran_bus):

\[x (\frac{ 4 }{ x })^2 =12\]

ganeshie8 (ganeshie8):

You must know what to do next

OpenStudy (amtran_bus):

Lol I should you are right but it sure does not mean I do Well, Can you say x 16/x^2 =12?

ganeshie8 (ganeshie8):

Keep going

OpenStudy (amtran_bus):

\[x(\frac{ 16 }{ x^{2} })=12\] and then you can multiply by x^2 \[x^{3} (16) =12?\]

ganeshie8 (ganeshie8):

Slow down your bus and relax

OpenStudy (amtran_bus):

The air compressor isn't working...no air in the air brakes.

OpenStudy (amtran_bus):

Well, If I am right so far, then 12/16

ganeshie8 (ganeshie8):

You cannot simply multiply a thing just one side. You must do the same thing to both sides.

OpenStudy (amtran_bus):

Shoot

OpenStudy (amtran_bus):

lol I feel like I'm at a math spa

ganeshie8 (ganeshie8):

You would be doing these problems quickly if you did them slowly and relaxedly

OpenStudy (amtran_bus):

Sorry ya'll it is so slow

OpenStudy (amtran_bus):

Ok so does the x^2 go on both sides?

OpenStudy (amtran_bus):

\[x^{3}(16)=12* x^{2}\]

OpenStudy (amtran_bus):

I guess he gone

ganeshie8 (ganeshie8):

Good. Keep going

OpenStudy (amtran_bus):

Oh.

OpenStudy (amtran_bus):

I don't understand why you had to multiply that on both sides. I mean, I was just simplifying the LHS

ganeshie8 (ganeshie8):

Oh wait, looks I've messed up. Let me pause my TV, one sec

ganeshie8 (ganeshie8):

We have x * 16/x^2 = 12

OpenStudy (amtran_bus):

Yes

ganeshie8 (ganeshie8):

Then you wanted to multiply both sides by x^2 is it ?

OpenStudy (amtran_bus):

x3(16)=12∗x2

ganeshie8 (ganeshie8):

No dear

OpenStudy (amtran_bus):

oh well I give up

ganeshie8 (ganeshie8):

On the left hand side, notice that x^2 is in the bottom

ganeshie8 (ganeshie8):

That's the reason you wanted to multiply x^2 both sides. You want to kill the denominator by multiplying both sides by it.

OpenStudy (amtran_bus):

I just posted that and you said no...

OpenStudy (amtran_bus):

@mathstudent55 please help

OpenStudy (agent0smith):

\[x *\frac{ 16 }{ x^2 } =12\]why not simplify first...\[\frac{ 16 }{ x }= 12\]